which number has the same property as the following numbers? 
(363,489,579) 
a)281   b)382   c)471   d)562

Grade:10

4 Answers

Latika Leekha
8 years ago
Hello student,
The numbers given to us are 363, 489, 579
By seeing these three numbers, we can judge that the sum of the digits in all three cases is 3 as,
363 = 3 + 6 + 3 = 12 = 1+2 = 3.
489 = 4 + 8 + 9 = 12 + 9 = 21 = 2 +1 = 3.
579 = 5 + 7 + 9 = 21 = 2 +1 = 3.
So from the options,
(a) 281 = 2 + 8 +1 = 11 = 1 + 1 = 2
(b) 382 = 3 + 8 + 2 = 13 = 1 + 3 = 4
(c) 471 = 4 + 7 + 1 = 12 = 1 + 2 = 3
(d) 562 = 5 + 6 + 2 = 13 = 1 + 3 = 4.
So, option (c) is correct.
grenade
8 years ago
the answer comes to be C because of the addition results in the number 3
Shubham Singh
8 years ago
the answer is c because it is the only multiple of 3
PRAPULPODISHETTI
8 years ago
The answer is 471 because all numbers 363,489,579 and our answer 471 are multiples of 3.